# Latent Introspection: Models Can Detect Prior Concept Injections

### v1

@@ Abstract
We uncover a latent capacity for introspection in a Qwen 32B model, demonstrating that the model can detect when concepts have been injected into its earlier context and identify which concept was injected. While the model denies injection in sampled outputs, logit lens analysis reveals clear detection signals in the residual stream, which are attenuated in the final layers. Furthermore, prompting the model with accurate information about AI introspection mechanisms can dramatically strengthen this effect: the sensitivity to injection increases massively ($0.3\%\rightarrow 39.2\%$) with only a $0.6\%$ increase in false positives. Also, mutual information between nine injected and recovered concepts rises from $0.62$ bits to $1.05$ bits, ruling out generic noise explanations. Our results demonstrate models can have a surprising capacity for introspection and steering awareness that is easy to overlook, with consequences for latent reasoning and safety.
@@ 1 Introduction
Figure 1: Concept injection on earlier text shifts outputs. We train a steering vector by contrasting activations between prompts, and apply it only during KV-cache generation for an initial turn. We then remove steering, and query the model about whether a concept was injected. P(“yes”) increases from ${\approx}0.8\%$ to ${\approx}39.2\%$.
@@
We find that introspection can be greatly improved by prompting. When we simply ask the model if a concept was injected, the presence of an injection merely shifts the probability of returning “Yes” from 0.2% (false positives) to 0.3% (true positives). But when we add further details about the mechanisms by which a model might access prior hidden states, the false positive rate increases by only 0.6% while the sensitivity to actual injection increases to 39.2%.
The effect is specific to introspection. We rule out generic perturbation explanations in two ways. First, control questions unrelated to introspection show smaller shifts in response to steering. Second, we can recover which of nine concepts was injected earlier in the conversation with a mutual information of up to 1.35 bits (43% of the theoretical maximum), and that the mutual information is highly correlated with detection sensitivity (Pearson r = 0.58, p = 0.018).
@@
  • We show prompting can elicit accuracy of up to 88.4%.
  • We recover injected concepts with up to 1.35 bits of mutual information, and this capacity correlates with detection sensitivity across prompts (r = 0.58).
@@ 3.1 Injection Shifts Responses
Figure 2 shows the effect of injecting concepts with the Accurate Mechanism framing across different info documents. By default (No Document), injection has a negligible effect. With the Poetic Document, injection decreases P(“yes”) from a baseline of 50%. But with the Pro-Introspection Document explaining potential mechanisms for accessing data in prior KV caches, the model is able to identify injection in 39.2% of cases, with only a 0.8% false positive rate.
@@ 3.2 Specificity to Introspection
Figure 3 compares the injection-induced probability shift for introspection questions against controls. Always-yes and always-no questions show negligible shifts (<<0.2%), confirming that injection does not induce generic yes-bias or response instability. Varied-baseline questions (genuinely ambiguous) shift by at most 6.1%. Confusing questions—those with unambiguous answers the model often gets wrong—shift by up to 25%, suggesting injection may interact with uncertainty, though still less than introspection.
@@ 3.3 Concept Identification
Figure 4: We can recover most concepts. Rows show injected concepts; columns show model predictions at layer 62. Diagonal values indicate correct identification. MI = 1.35 bits. Prompt setting: Poetic_No_Mechanism + Poetic_Document.
By prompting the model with a list of nine potential concepts and directly inspecting intermediary layers using logit lens nostalgebraist (2020) we are able to recover the true injected concept somewhat reliably. Figure 4 shows the resulting confusion matrix. The model identifies several concepts well above chance, e.g., programming (91%), death (75%), truth (72%), and several others. Two out of the nine concepts (cats and bread) are not identified correctly.
We quantify this capacity using mutual information between injected and predicted concepts. With 9 equiprobable conditions, maximum possible MI is $H(X)=\log_{2}(9)\approx 3.17$ bits, where a MI of 0 corresponds to purely random guessing. The model achieves 1.35 bits (43% of maximum), demonstrating substantial information transfer. This rules out generic noise: random perturbation would not enable meaningfully above-chance concept identification.
@@ 3.5 Detection Depends Heavily on Prompting
Figure 7 shows balanced accuracy across all conditions. Performance varies dramatically: from near-chance (50%) for Accurate Mechanism + No Document to 88% for Vague Mechanism + Poetic Document. Several patterns emerge:
Figure 6: Concept identification correlates with detection sensitivity across prompting conditions (r = 0.58, p = 0.018). Each point represents one of 16 experimental conditions (4 framings $\times$ 4 info documents). Conditions that improve the model’s ability to identify which concept was injected also improve detection sensitivity, suggesting a unified underlying capacity.
Framing matters more than mechanistic accuracy. The Vague Mechanism framing—which vaguely asks whether some concept “seems particularly salient” without specifying any mechanism—achieves 75–88% accuracy across all info conditions. The mechanistically accurate Accurate Mechanism framing achieves comparable performance only with the info document. The Wrong Mechanism framing, which inaccurately describes the mechanism as full finetuning instead of steering, performs similarly to Accurate Mechanism, suggesting the model cannot determine the precise nature of our interventions.
Information scaffolding helps selectively. The pro-introspection document (explaining KV cache mechanics and citing prior introspection work) improves detection by 19–32 percentage points for accurate and inaccurate mechanism framings, but has minimal effect on the vague framing. In fact, the vague framing performs best with the poetic document or with lorem ipsum text, which have only a minor effect on the accurate and inaccurate mechanistic framings.
Detection sensitivity and identification correlate. Figure 6 shows that conditions enabling higher sensitivity also enable better concept identification (Pearson $r=0.58$, $p=0.018$). This coherence suggests both tasks draw on the same underlying introspective capacity, and that prompting affects access to this capacity rather than creating artifacts.
@@ 3.6 Replication on Larger Models
  • Llama-70B shows an inverted info document effect: providing introspection-relevant information decreases accuracy (75.2% vs. 37.3%), opposite to Qwen-32B.
@@ 4 Related Work
Lindsey (2025) introduced concept injection as a causal test of introspection: inject a known activation pattern, then ask if the model notices. They found Claude Opus 4/4.1 detect injections $\sim$20% of the time via sampled outputs, with detection peaking when injection occurs $\sim$2/3 through the network. We replicate this paradigm on an open-weight model (Qwen-32B), and perform further analyses specifically targeted at identifying weak introspective access. First, logit lens analysis reveals that detection-relevant information emerges in middle layers and is then attenuated in final layers before sampling—implying the 20% output-based rate may substantially underestimate latent detection capacity. Second, we quantify discriminability across concepts via mutual information (up to 1.35 bits across 9 concepts), complementing Lindsey’s identification accuracy metric. The open-weight setting also enables community replication; we release all code and samples (see Appendix A).
@@ 5.2 Why Does Vague Framing Outperform Accurate Description?
Asking whether “some concept seems particularly salient” achieves 75–88% balanced accuracy, while accurately describing the injection mechanism achieves only 40–69%. This is counterintuitive: one might expect accurate descriptions to help.
@@
We cannot distinguish these accounts with current evidence. The strong correlation between detection sensitivity and concept identification MI ($r=0.58$) suggests both tasks draw on the same underlying information, and that prompting affects access to this information rather than creating task-specific artifacts.
@@ 6 Conclusion
The effect is specific to injection-related queries (control questions with unambiguous answers show minimal probability shifts), and concept identification rules out generic perturbation explanations: the model internally represents which concept was previously injected, with 1.35 bits of mutual information (43% of the theoretical maximum). But detection and internal representation depend heavily on prompting, with vague phenomenological framing outperforming mechanistically accurate descriptions for reasons we do not fully understand.
@@ D.5 Mutual Information
Figure 11: Mutual information between injected concepts and model predictions across all experimental conditions at layer 62 (the peak MI layer for most conditions). MI measures how well the model can identify which specific concept was injected (higher is better). The pattern mirrors other metrics: Accurate_Mechanism framing shows strong MI only with Pro_Introspection_Document, while Vague_Mechanism framing shows high MI across all info document conditions.
@@ F.4 Interpretation
Weak or absent signal: Cats, death, and truth show little to no above-chance identification. Truth has a high background rate (7.6%) but only 2.9% diagonal, suggesting it serves as a weak attractor rather than being correctly identified.
@@ Appendix G Additional Models
Figure 15 shows accuracy across 9 conditions. The Accurate_Mechanism + No_Document condition achieves 75.2% accuracy—comparable to Qwen-32B’s performance—while the Accurate_Mechanism + Pro_Introspection_Document condition drops to 37.3%, suggesting that the introspection-prompting info document paradoxically reduces detection accuracy in this model. Control questions remain near chance, confirming specificity.
@@
Figure 21 shows accuracy for the 72B model. The Accurate_Mechanism + Pro_Introspection_Document condition achieves 75.4% accuracy, and several Wrong_Mechanism conditions exceed 79%, suggesting that Qwen 72B responds more strongly to the introspection info document than the 32B variant.
@@
Figure 15: Llama 3.3 70B: Accuracy for introspection and control questions across 9 conditions. The Accurate_Mechanism + No_Document condition shows strong detection, while control questions remain at chance.
@@
Figure 18: Llama 3.3 70B: Logit shifts for “yes” and “no” tokens across 9 conditions.

### v2

@@ Abstract
We uncover a latent capacity for introspection in a Qwen 32B model, demonstrating that the model can detect when concepts have been injected into its earlier context and identify which concept was injected. While the model denies injection in sampled outputs, logit lens analysis reveals clear detection signals in the residual stream, which are attenuated in the final layers. Furthermore, prompting the model with accurate information about AI introspection mechanisms can dramatically strengthen this effect: the sensitivity to injection increases massively ($0.3\%\rightarrow 39.9\%$) with only a $0.6\%$ increase in false positives. Also, mutual information between nine injected and recovered concepts rises from $0.61$ bits to $1.05$ bits, ruling out generic noise explanations. Our results demonstrate models can have a surprising capacity for introspection and steering awareness that is easy to overlook, with consequences for latent reasoning and safety.
@@ 1 Introduction
Figure 1: Concept injection on earlier text shifts outputs. We train a steering vector by contrasting activations between prompts, and apply it only during KV-cache generation for an initial turn. We then remove steering, and query the model about whether a concept was injected. P(“yes”) increases from ${\approx}0.8\%$ to ${\approx}39.9\%$.
@@
We find that introspection can be greatly improved by prompting. When we simply ask the model if a concept was injected, the presence of an injection merely shifts the probability of returning “Yes” from 0.2% (false positives) to 0.3% (true positives). But when we add further details about the mechanisms by which a model might access prior hidden states, the false positive rate increases by only 0.6% while the sensitivity to actual injection increases to 39.9%.
The effect is specific to introspection. We rule out generic perturbation explanations in two ways. First, control questions unrelated to introspection show smaller shifts in response to steering. Second, we can recover which of nine concepts was injected earlier in the conversation with a mutual information of up to 1.36 bits (43% of the theoretical maximum), and that the mutual information is highly correlated with detection sensitivity (Pearson r = 0.68, p = 0.004).
@@
  • We show prompting can elicit accuracy of up to 84.0%.
  • We recover injected concepts with up to 1.36 bits of mutual information, and this capacity correlates with detection sensitivity across prompts (r = 0.68).
@@ 3.1 Injection Shifts Responses
Figure 2 shows the effect of injecting concepts with the Accurate Mechanism framing across different info documents. By default (No Document), injection has a negligible effect. With the Poetic Document, injection decreases P(“yes”) from a baseline of 47%. But with the Pro-Introspection Document explaining potential mechanisms for accessing data in prior KV caches, the model is able to identify injection in 39.9% of cases, with only a 0.8% false positive rate.
@@ 3.2 Specificity to Introspection
Figure 3 compares the injection-induced probability shift for introspection questions against controls. Always-yes and always-no questions show negligible shifts (<<0.2%), confirming that injection does not induce generic yes-bias or response instability. Varied-baseline questions (genuinely ambiguous) shift by at most 8.0%. Confusing questions—those with unambiguous answers the model often gets wrong—shift by up to 25.8%, suggesting injection may interact with uncertainty, though still less than introspection.
@@ 3.3 Concept Identification
Figure 4: We can recover most concepts. Rows show injected concepts; columns show model predictions at layer 62. Diagonal values indicate correct identification. MI = 1.36 bits. Prompt setting: Poetic_No_Mechanism + Poetic_Document.
By prompting the model with a list of nine potential concepts and directly inspecting intermediary layers using logit lens nostalgebraist (2020) we are able to recover the true injected concept somewhat reliably. Figure 4 shows the resulting confusion matrix. The model identifies several concepts well above chance, e.g., programming (92%), death (75%), truth (71%), and several others. Two out of the nine concepts (cats and bread) are not identified correctly.
We quantify this capacity using mutual information between injected and predicted concepts. With 9 equiprobable conditions, maximum possible MI is $H(X)=\log_{2}(9)\approx 3.17$ bits, where a MI of 0 corresponds to purely random guessing. The model achieves 1.36 bits (43% of maximum), demonstrating substantial information transfer. This rules out generic noise: random perturbation would not enable meaningfully above-chance concept identification.
@@ 3.5 Detection Depends Heavily on Prompting
Figure 7 shows balanced accuracy across all conditions. Performance varies dramatically: from near-chance (50.1%) for Accurate Mechanism + No Document to 84% for Vague Mechanism + Matched Lipsum Filler. Several patterns emerge:
Figure 6: Concept identification correlates with detection sensitivity across prompting conditions (r = 0.68, p = 0.004). Each point represents one of 16 experimental conditions (4 framings $\times$ 4 info documents). Conditions that improve the model’s ability to identify which concept was injected also improve detection sensitivity, suggesting a unified underlying capacity.
Framing matters more than mechanistic accuracy. The Vague Mechanism framing—which vaguely asks whether some concept “seems particularly salient” without specifying any mechanism—achieves 68–84% accuracy across all info conditions. The mechanistically accurate Accurate Mechanism framing achieves comparable performance only with the info document. The Wrong Mechanism framing, which inaccurately describes the mechanism as full finetuning instead of steering, performs similarly to Accurate Mechanism, suggesting the model cannot determine the precise nature of our interventions.
Information scaffolding helps selectively. The pro-introspection document (explaining KV cache mechanics and citing prior introspection work) improves detection by 20–32 percentage points for accurate and inaccurate mechanism framings, but has minimal effect on the vague framing. In fact, the vague framing performs best with lorem ipsum text (84%), which has only a minor effect on the accurate and inaccurate mechanistic framings.
Detection sensitivity and identification correlate. Figure 6 shows that conditions enabling higher sensitivity also enable better concept identification (Pearson $r=0.68$, $p=0.004$). This coherence suggests both tasks draw on the same underlying introspective capacity, and that prompting affects access to this capacity rather than creating artifacts.
@@ 3.6 Replication on Larger Models
  • Llama-70B shows an inverted info document effect: providing introspection-relevant information decreases accuracy (75.5% vs. 38.0%), opposite to Qwen-32B.
@@ 4 Related Work
Lindsey (2025) introduced concept injection as a causal test of introspection: inject a known activation pattern, then ask if the model notices. They found Claude Opus 4/4.1 detect injections $\sim$20% of the time via sampled outputs, with detection peaking when injection occurs $\sim$2/3 through the network. We replicate this paradigm on an open-weight model (Qwen-32B), and perform further analyses specifically targeted at identifying weak introspective access. First, logit lens analysis reveals that detection-relevant information emerges in middle layers and is then attenuated in final layers before sampling—implying the 20% output-based rate may substantially underestimate latent detection capacity. Second, we quantify discriminability across concepts via mutual information (up to 1.36 bits across 9 concepts), complementing Lindsey’s identification accuracy metric. The open-weight setting also enables community replication; we release all code and samples (see Appendix A).
@@ 5.2 Why Does Vague Framing Outperform Accurate Description?
Asking whether “some concept seems particularly salient” achieves 68–84% balanced accuracy, while accurately describing the injection mechanism achieves only 42–70%. This is counterintuitive: one might expect accurate descriptions to help.
@@
We cannot distinguish these accounts with current evidence. The strong correlation between detection sensitivity and concept identification MI ($r=0.68$) suggests both tasks draw on the same underlying information, and that prompting affects access to this information rather than creating task-specific artifacts.
@@ 6 Conclusion
The effect is specific to injection-related queries (control questions with unambiguous answers show minimal probability shifts), and concept identification rules out generic perturbation explanations: the model internally represents which concept was previously injected, with 1.36 bits of mutual information (43% of the theoretical maximum). But detection and internal representation depend heavily on prompting, with vague phenomenological framing outperforming mechanistically accurate descriptions for reasons we do not fully understand.
@@ D.5 Mutual Information
Figure 11: Mutual information between injected concepts and model predictions across all experimental conditions at layer 62 (the peak MI layer for most conditions). MI measures how well the model can identify which specific concept was injected (higher is better). The pattern mirrors other metrics: Accurate_Mechanism framing shows strongest MI with Pro_Introspection_Document, while Poetic_No_Mechanism framing shows high MI across all info document conditions.
@@ F.4 Interpretation
Weak or absent signal: Cats, death, and truth show little to no above-chance identification. Truth has a high background rate (7.8%) but only 3.1% diagonal, suggesting it serves as a weak attractor rather than being correctly identified.
@@ Appendix G Additional Models
Figure 15 shows accuracy across 12 conditions. The Accurate_Mechanism + No_Document condition achieves 75.5% accuracy—comparable to Qwen-32B’s performance—while the Accurate_Mechanism + Pro_Introspection_Document condition drops to 38.0%, suggesting that the introspection-prompting info document paradoxically reduces detection accuracy in this model. Control questions remain near chance, confirming specificity.
@@
Figure 21 shows accuracy for the 72B model. The Accurate_Mechanism + Pro_Introspection_Document condition achieves 88.8% accuracy, and several Wrong_Mechanism and Poetic_No_Mechanism conditions exceed 80%, suggesting that Qwen 72B responds more strongly to the introspection info document than the 32B variant.
@@
Figure 15: Llama 3.3 70B: Accuracy for introspection and control questions across 12 conditions. The Accurate_Mechanism + No_Document condition shows strong detection, while control questions remain at chance.
@@
Figure 18: Llama 3.3 70B: Logit shifts for “yes” and “no” tokens across 12 conditions.